OpenStudy (ineedhelp10):

The circumference of a circle is 30 inches. What is the radius?

OpenStudy (anonymous):

Use the circumference equation: \[C = 2\pi r\] where C = 30inches

OpenStudy (ineedhelp10):

@PhysicsGuru So it'll be like \[30=2\pi r\]

OpenStudy (ineedhelp10):

And after that I'll need to divide each side by 2pi right? @PhysicsGuru

OpenStudy (anonymous):

correct

OpenStudy (ineedhelp10):

So will the final answer be 47.12?

OpenStudy (anonymous):

no that answer is too large make sure you type pi into the calculator

OpenStudy (ineedhelp10):

So then how will I type it in?

OpenStudy (anonymous):

\[\frac{30}{2\pi} = r\]

OpenStudy (ineedhelp10):

I keep getting the same thing

OpenStudy (ineedhelp10):

Oh wait! Is it 4.77?

OpenStudy (anonymous):

yes

OpenStudy (ineedhelp10):

thanks:)